OpenStudy (anonymous):

There are 3 Xbox games and 5 Wii games left on the shelf. What is the ratio of Xbox games to Wii games? 3:5 3:8 5:3

OpenStudy (anonymous):

3:5

OpenStudy (anonymous):

could u help me with 2 more questions?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

What is the value of X? 7/x = 3/15 21 14 35

OpenStudy (anonymous):

the answer is 35

OpenStudy (anonymous):

Convert the following decimal to a percent! 0.35 0.35% 3.5% 35%

OpenStudy (anonymous):

7/x = 3/15 cross-multiply: 3x=105 105/3=35

OpenStudy (anonymous):

the answer is 35%

OpenStudy (anonymous):

yea 35%

OpenStudy (anonymous):

omg thank u so much!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:) yw